38) Which of the following occurs naturally as nonpolar diatomic molecules? 

A) sulfur B) argon C) chlorine D) all of the above E) none of the above 


39) Aspartic acid is an amino acid used to synthesize proteins. How many lonc pairs of electrons need to be added to complete the aspartic acid structure shown below? 

 A) 11 B) 9 096 D) 8 E) 4 


40) The number of resonance structures for the sulfur dioxide molecule that satisfy the octet rule is 

A) 4 B) 3 C) 1 D) 2 E) None of these
Answer #1

Q38. Correct answer is : (C) chlorine

Q39. Correct answer is : (B) 9

Q40. Correct answer is : (D) 2

Explanation

Q38. Sulfur naturally occurs as S8 molecule whereas argon occurs as a monoatomic gas Ar.

Only chlorine occurs as nonpolar diatomic molecule Cl2

Q39. In the given aspartic acid structure, there are 4 oxygen atoms. Each oxygen atom requires 2 lone pairs of electrons. Hence, total lone pairs for oxygen are 4 * 2 = 8

Also, there is one nitrogen atom which requires one lone pair of electron. Total lone pairs for nitrogen are 1 * 1 = 1

Total lone pairs = 8 + 1 = 9